# SLogic: Subgraph-Informed Logical Rule Learning for Knowledge Graph Completion

### v1

@@ Abstract
Logical rule-based methods offer an interpretable approach to knowledge graph completion by capturing compositional relationships in the form of human-readable inference rules. However, current approaches typically treat logical rules as universal, assigning each rule a fixed confidence score that ignores query-specific context. This is a significant limitation, as a rule’s importance can vary depending on the query. To address this, we introduce SLogic (Subgraph-Informed Logical Rule learning), a novel framework that assigns query-dependent scores to logical rules. The core of SLogic is a scoring function that utilizes the subgraph centered on a query’s head entity, allowing the significance of each rule to be assessed dynamically. Extensive experiments on benchmark datasets show that by leveraging local subgraph context, SLogic consistently outperforms state-of-the-art baselines, including both embedding-based and rule-based methods.
@@ 1 Introduction
Knowledge graphs (KGs) capture complex relationships among diverse real-world entities, including commonly used ones such as people, locations, organizations, events, products, and concepts, as well as more specialized types such as genes, diseases, chemicals, publications, and technical terms. They are foundational to many applications, such as recommendation systems and question answering Hildebrandt et al. (2019); Lan & Jiang (2020). For instance, in a movie KG, a recommendation system can suggest a new film to a user by identifying that they have liked other movies directed by the same person or starring a particular actor.
@@
Despite their usefulness, KGs are notoriously incomplete. To address this, the field of Knowledge Graph Completion (KGC) has explored several dominant paradigms. Embedding-based models (e.g., (Bordes et al., 2013; Yang et al., 2014; Dettmers et al., 2018; Trouillon et al., 2016; Sun et al., 2019)) learn vector representations of entities and relations to predict missing links. More recently, Graph Neural Networks (GNNs) based models (Zhang & Yao, 2022; Zhu et al., 2021; 2023) have achieved state-of-the-art performance by capturing intricate topological patterns within the graph’s structure. However, the strength of both these approaches is also their primary weakness: their reasoning is opaque. Because they operate on dense, sub-symbolic vectors, they function as “black box” models, making it difficult to understand or trust their predictions in critical applications.
@@
As an alternative, rule-based reasoning offers a transparent and explainable approach to KGC. These methods infer missing links by discovering and applying logical rules (e.g., $bornIn(X,Y)\land locatedIn(Y,Z)\rightarrow livesIn(X,Z)$), which are inherently human-readable. This interpretability is a significant advantage, providing not just a prediction, but also the logical path to reach it. While rule-based models offer inductive capabilities and can outperform many embedding-based methods, a common limitation is that they treat rules as universally static, assigning a fixed confidence score to each rule regardless of the context. This overlooks the fact that a rule’s relevance can change dramatically depending on the specific query, and therefore, answers to a query should be context-dependent.
@@
For a concrete example, consider the KGC query $livesIn(Person,?)$. One plausible rule, $bornIn(X,Y)\land locatedIn(Y,Z)\rightarrow livesIn(X,Z)$, might link a person to their birth location. A second rule, $worksAt(X,Y)\land locatedIn(Y,Z)\rightarrow livesIn(X,Z)$, could link them to their place of work. A model relying on static rule confidences, perhaps favoring the globally more common “birthplace” rule, would fail to distinguish the second rule is more relevant in the context of a CEO of a major tech company because the “worksAtworksAt” path becomes exceptionally reliable. Our approach uses query-specific subgraphs to dynamically assess the relevance of each potential reasoning path for a given query, capturing the nuance that static models miss.
To capture and utilize the local importance of a rule, we propose a novel hybrid approach that enhances rule-based reasoning with contextual awareness. Our key innovation is to generate query-specific rule weights by synergizing the strengths of both symbolic rules and neural networks. We first pre-calculate simple paths to serve as our explicit, interpretable rule base. We then employ a GNN not as the final predictor, but as a powerful context encoder for the query. By integrating the GNN-derived context with the logical paths, our model can dynamically assess which rules are most significant for any given query. This method retains the core explainability of a rule-based system while leveraging the contextual power of GNNs to achieve more accurate and nuanced inferences.
@@ 2 Related Work
The literature on knowledge graph completion can be broadly categorized into embedding-based approaches (Bordes et al., 2013; Yang et al., 2014; Dettmers et al., 2018; Trouillon et al., 2016; Sun et al., 2019), graph neural network based methods (Zhang & Yao, 2022; Zhu et al., 2021; 2023), and rule-based strategies. As explained in Section 1, rule-based reasoning enhances the explainability of knowledge graph completion. We therefore focus on describing such methods.
Existing rule-based methods can be broadly classified according to how they discover and learn logical rules. A line of work focuses on learning rule confidences by actively searching for paths (groundings) on the graph during the training process. NeuralLP (Yang et al., 2017) and its successor DRUM (Sadeghian et al., 2019) use a differentiable framework inspired by TensorLog (Cohen et al., 2017) to find soft proofs. Similarly, RNNLogic (Qu et al., 2021) is a probabilistic model that treats logic rules as a latent variable and uses an Expectation-Maximization (EM) algorithm to simultaneously train a rule generator and a reasoning predictor. The second paradigm defers the search for rules at inference time, aiming to discover novel patterns not explicitly seen during training. RLogic (Cheng et al., 2022) and NCRL (Cheng et al., 2023) are recent examples that learn a neural function to score the quality of potential rule structures. They perform an exhaustive, brute-force enumeration of all possible rules up to a certain length. All prior efforts rely on rules deemed important at the level of the entire knowledge graph. In contrast, our proposed SLogic framework recalculates rule importance with respect to contextual entities and their surrounding subgraphs.
@@ 4 The SLogic Framework
Unlike traditional methods that assign static confidence scores to logical rules, SLogic learns a dynamic, query-dependent scoring function $\phi(h,r,\mathbf{r}_{b})$. This function assesses the relevance of a candidate rule $r\leftarrow\mathbf{r}_{b}$ by considering not only the rule itself but also the rich structural context of the query’s head entity, $h$. In order to calculate a context-aware score for a relation $r$, we design a novel strategy to generate instance that are enriched by context-aware rules and to train a model.
@@ 4.1.1 Rule Base Construction
The initial step of our method is to mine a comprehensive set of logical rules from the training knowledge graph $\mathcal{G}$. To extract these rules, we iterate through each ground truth triple $(h,r_{h},t)$ in the training set. For each triple, we treat it as a positive example of a rule instantiation where $r_{h}$ is the head relation. We then perform a Depth-First Search (DFS) to find all simple paths connecting the head entity $h$ and the tail entity $t$ within the graph, up to the predefined maximum length $L$. Each discovered path is converted into its corresponding relational path, which forms a candidate rule body $\mathbf{r}_{b}$.
After enumerating all candidate rules, we quantify their quality using established statistical metrics. For each rule $r_{h}\leftarrow\mathbf{r}_{b}$, we compute its standard confidence, defined as the conditional probability $P(r_{h}|\mathbf{r}_{b})$,
@@ 4.1.2 Query-Centric Subgraph Extraction
Feature engineering. A raw subgraph is insufficient for a GNN to understand the relative importance of its components. We therefore enrich the subgraph by engineering a set of node features, $\mathbf{x}_{v}$, for each node $v\in\mathcal{G}_{h}$. These features are designed to encode the structural role of each node relative to the central query entity $h$. The feature vector $\mathbf{x}_{v}$ is a concatenation of three components. (i) Head Entity Indicator: A two-dimensional binary vector that identifies whether a node is the query head entity $h$ (i.e., $[1,0]$) or a neighbor (i.e., $[0,1]$). This anchors the subgraph representation around the query. (ii) Shortest Path Distance: The geodesic distance (i.e., minimum number of hops) from node $v$ to the head entity $h$ within the subgraph $\mathcal{G}_{h}$. This explicitly encodes the structural proximity of each node to the query’s origin. (iii) Global Centrality Score: The log-scaled degree of node $v$ as calculated from the complete graph $\mathcal{G}$. This feature injects a measure of the node’s global importance into the local subgraph context.
@@ 4.1.3 Instance creation and Negative Sampling
For a query $(h_{i},r_{i},t_{i})$, let ${\bf R}_{i}^{+}$ and ${\bf R}_{i}^{-}$ consist of the $k_{pos}$ positive rules $r_{i}\leftarrow r_{b_{i1}}^{+},\cdots,r_{i}\leftarrow r_{b_{ik_{pos}}}^{+}$} and $k_{neg}$ hard negative rules $r_{i}\leftarrow r_{b_{i1}}^{-},\cdots,r_{i}^{-}\leftarrow r_{b_{ik_{neg}}}^{-}$} respectively. The instances generated in this step, subsequently provided to our model, comprise rule-enriched triplets, $((h_{1},r_{1},t_{1}),{\bf R}_{1}^{+}$, ${\bf R}_{i}^{-})$, $((h_{2},r_{2},t_{2}),{\bf R}_{2}^{+}$, ${\bf R}_{2}^{-})$, etc.
@@ 4.3 Inference and Final Ranking
Input: query $\mathbf{q}(h,r,?)$, model $\phi(\cdot)$, hyperparameters Output: a vector of scores that measure the plausibility of different entities answering $\mathbf{q}$ 1. Candidate rule generation from global rule base 2. For each rule $r\leftarrow\mathbf{r}_{b}$, predict $\phi(h,r,\mathbf{r}_{b})$ from model $\phi(\cdot)$ 3. Update each rule’s score $\phi$ to a confidence score $w_{i}$ using softmax with temperature 4. For all the entities, calculate their potential to answer $\mathbf{q}$ (a) Calculate a rule grounding score $ground(h,\mathbf{r}_{b})$ (b) Apply a binarization function to the grounding score $B(ground(h,\mathbf{r}_{b}))$ (c) Compute $\mathbf{v}_{\text{ans}}$ using equation 2, which represents the entities potential to answer $\mathbf{q}$ 5. Return $\mathbf{v}_{\text{ans}}$
@@
Figure 2 shows the steps to answer the given query.
Step 1 generates candidate rules. For any given query $(h,r,?)$, exhaustively scoring all rules in the global rule base would be computationally prohibitive. To obtain a manageable yet high-quality candidate set, we first identify all rules that are locally applicable to $h$, and then select the top-$N$ among them based on their Wilson score.
Step 2 calculates a context-aware (localized) importance score for each candidate rule. In particular, each candidate rule $\mathbf{r}_{b_{i}}$ is passed to our trained scoring function $\phi(h,r,\mathbf{r}_{b_{i}})$, which leverages the query’s unique subgraph context $\mathcal{G}_{h}$. This step assigns a dynamic, context-aware score to each rule, moving beyond static metrics to assess its true relevance for the specific query. Next, the context-aware scores are converted into a confidence distribution using a temperature-controlled softmax (Step 3). The weight $w_{i}$ for each rule is calculated as $w_{i}=\frac{\exp(\phi^{\prime}_{i}/T)}{\sum_{j}\exp(\phi^{\prime}_{j}/T)}$. The temperature $T$ is an important hyperparameter. As $T\to 0$, the softmax approximates an argmax function, concentrating all weight onto the single highest-scoring rule. This allows the model to rely solely on the strongest contextual signal.
@@
A naive grounding score, can be a noisy signal dominated by high-degree nodes. To mitigate this effect, we squash the path counts using the tanh function, controlled by a tanh scale hyperparameter $\tau$, $B(\mathbf{g}_{i})=\tanh(\mathbf{g}_{i}/\tau)$. This provides a soft, saturated count that is controlled by the tanh scale hyperparameter $\tau$, preventing an unbounded influence from numerous paths.
Finally (Step 4c), a vector $\mathbf{v}_{\text{ans}}$ that captures all the entities’ plausibility to answer the given query is computed as the weighted sum of all the binarized, grounded rule vectors as in equation 2:
@@ 5 Experiments: SLogic’s performance in KGC
Default setting of hyperparameters. For rule base construction, the length of rule body (or the depth of the DFS) $L$ is set to 5 (for WN18RR) and 3 (for both FB15K-237 and YAGO3-10). In subgraph extraction, the number of subgraph hops $k$ is set to be 4 for WN18RR dataset and 1 for both FB15K-237 and YAGO3-10 datasets. The hyperparameter $\alpha$ is 100 for all the datasets. For model training, the margin $\epsilon$ is set to be 1, $k_{pos}$ and $k_{neg}$ are set to be 5 and 20 respectively for all the three datasets. During inference stage, we set the default parameters as $N=50$ for WN18RR and FB15k-237, and $N=10$ for YAGO3-10. For all datasets, we use $T=0.5$ and $\tau=2.0$.
Further details on the experimental setting, including hardware configuration, implementation specifics, and training procedures, are provided in Appendix B.

### v2

@@ Abstract
Logical rule-based methods offer an interpretable approach to knowledge graph completion (KGC) by capturing compositional relationships in the form of human-readable inference rules. While existing logical rule-based methods learn rule confidence scores, they typically assign a global weight to each rule schema, applied uniformly across the graph. This is a significant limitation, as a rule’s importance often varies depending on the specific query instance. To address this, we introduce SLogic (Subgraph-Informed Logical Rule learning), a novel framework that assigns query-dependent scores to logical rules. The core of SLogic is a context-aware scoring function. This function determines the importance of a rule by analyzing the subgraph locally defined by the query’s head entity, thereby enabling a differentiated weighting of rules specific to their local query contexts. Extensive experiments on benchmark datasets show that SLogic outperforms existing rule-based methods and achieves competitive performance against state-of-the-art baselines. It also generates query-dependent, human-readable logical rules that serve as explicit explanations for its inferences.
@@ 1 Introduction
Knowledge graphs (KGs) capture complex relationships among diverse real-world entities, including commonly used ones such as people, locations, organizations, events, products, and concepts, as well as more specialized types such as genes, diseases, chemicals, publications, and technical terms. They are foundational to many applications, such as recommendation systems and question answering Hildebrandt et al. (2019); Lan & Jiang (2020).
@@
Despite their usefulness, KGs are notoriously incomplete. To address this, the field of Knowledge Graph Completion (KGC) has explored several dominant paradigms. Embedding-based models (e.g., (Bordes et al., 2013; Yang et al., 2014; Dettmers et al., 2018; Trouillon et al., 2016; Sun et al., 2019)) learn vector representations of entities and relations to predict missing links. More recently, Graph Neural Networks (GNNs) based models (Zhang & Yao, 2022; Zhu et al., 2021; 2023) have achieved state-of-the-art performance by capturing intricate topological patterns within the graph’s structure. However, the strength of both these approaches is also their primary weakness: their reasoning is opaque. They operate on dense, sub-symbolic vectors, they function as “black box” models, making it difficult to understand or trust their predictions in critical applications.
@@
As an alternative, rule-based reasoning offers a transparent and explainable approach to KGC. These methods infer missing links by discovering and applying logical rules (e.g., $bornIn(X,Y)\land locatedIn(Y,Z)\rightarrow livesIn(X,Z)$), which are inherently human-readable. This interpretability is a significant advantage, providing not just a prediction, but also the logical path to reach it. While rule-based models offer inductive capabilities and can outperform many embedding-based methods, a common limitation is that they learn one weight for each rule across the entire graph, despite the fact that weight learning takes into account the grounding context (e.g., (Yang et al., 2017; Sadeghian et al., 2019)). This overlooks the fact that a rule’s relevance can change depending on the specific query instance (e.g., a rule may be 80% confident for one query but only 50% for another), and therefore, answers to a query should be context-dependent.
@@
For a concrete example, consider the KGC query $livesIn(Person,?)$. One plausible rule, $bornIn(X,Y)\land locatedIn(Y,Z)\rightarrow livesIn(X,Z)$, might link a person to their birth location. A second rule, $worksAt(X,Y)\land locatedIn(Y,Z)\rightarrow livesIn(X,Z)$, could link them to their place of work. A model relying on global rule confidences, perhaps favoring the globally more common “birthplace” rule, would fail to distinguish the second rule is more relevant in the context of a CEO of a major tech company because the “worksAtworksAt” path becomes exceptionally reliable. Our approach uses query-specific subgraphs to dynamically assess the relevance of each potential reasoning path for a given query, capturing the nuance that global scoring models miss.
To capture and utilize the local importance of a rule, we propose a novel hybrid approach, SLogic (Subgraph-Informed Logical Rule learning), to enhance rule-based reasoning with contextual awareness. SLogic pre-calculates simple paths to serve as then explicit, interpretable rule base. It then employs a GNN not as the final predictor, but as a powerful context encoder for the query. The contribution of SLogic is twofold. (i) It learns query-dependent weights that improve prediction accuracy, particularly on datasets where global rules are ambiguous. (ii) The learned scores provide granular insight into the reasoning process, identifying not just which rules are invoked, but how much each contributes to completing a specific edge. To the best of our knowledge, SLogic is the first neuro-symbolic framework to move beyond learning global rule confidence, introducing a mechanism to learn rule weights tailored to specific query contexts. We emphasize that while SLogic outperforms other rule-based approaches, it is best understood as an interpretable rule-learning framework rather than a purely metric-driven link predictor.
@@ 2 Related Work
The literature on knowledge graph completion (KGC) can be broadly categorized into embedding-based approaches (Bordes et al., 2013; Yang et al., 2014; Dettmers et al., 2018; Trouillon et al., 2016; Sun et al., 2019), GNN-based methods (Zhang & Yao, 2022; Zhu et al., 2021; 2023), and rule-based strategies. Among GNN-based methods, NBFNet (Zhu et al., 2021) has achieved state-of-the-art performance by capturing all-path information through a neural Bellman-Ford message passing mechanism. While such methods are highly effective, they generally either operate as black boxes or provide path weights without learning explicit, human-readable rules that rule-based approaches offer. As explained in Section 1, rule-based reasoning enhances the explainability of knowledge graph completion. We therefore focus on describing such methods.
Existing rule-based methods can be broadly classified according to how they discover and learn logical rules. Probabilistic rule learning frameworks, such as Markov Logic Networks (MLNs) (Richardson & Domingos, 2006) and pLogicNet (Qu & Tang, 2019), combine logic with probabilistic models to learn weights for logical rules. A different line of work focuses on learning rule confidences by actively searching for paths (groundings) on the graph during the training process. NeuralLP (Yang et al., 2017) and its successor DRUM (Sadeghian et al., 2019) use a differentiable framework inspired by TensorLog (Cohen et al., 2017) to find soft proofs. Similarly, RNNLogic (Qu et al., 2021) is a probabilistic model that treats logic rules as a latent variable and uses an Expectation-Maximization (EM) algorithm to simultaneously train a rule generator and a reasoning predictor. Recent approaches such as RLogic (Cheng et al., 2022) and NCRL (Cheng et al., 2023) learn a neural function to score the quality of potential rule structures. They perform an exhaustive, brute-force enumeration of all possible rules up to a certain length. Another line of research explores hybrid frameworks that jointly learn embeddings and logical rules. Methods such as IterE (Zhang et al., 2019) and RPJE (Niu et al., 2020) iteratively update both components, where learned rules help refine entity and relation embeddings, and embeddings in turn guide the rule mining process. These prior efforts learn or rely on rules deemed important at the level of the entire knowledge graph. In contrast, our SLogic framework recalculates rule importance with respect to contextual entities and their surrounding subgraphs.
@@ 4 The SLogic Framework
Unlike traditional methods that assign static confidence scores to logical rules, SLogic learns a dynamic, query-dependent scoring function $\phi(h,r,\mathbf{r}_{b})$. This function assesses the relevance of a candidate rule $r\leftarrow\mathbf{r}_{b}$ by considering not only the rule itself but also the rich structural context of the query’s head entity, $h$. In order to calculate a context-aware score for a relation $r$, we design a novel strategy to generate instances that are enriched by context-aware rules and to train a model.
@@ 4.1.1 Rule Base Construction
The initial step of our method is to mine a comprehensive set of logical rules from the training knowledge graph $\mathcal{G}$. To extract these rules, we iterate through each ground truth triple $(h,r_{h},t)$ in the training set. For each triple, we treat it as a positive example of a rule instantiation where $r_{h}$ is the head relation. We then perform a Depth-First Search (DFS) to find all simple paths connecting the head entity $h$ and the tail entity $t$ within the graph, up to the predefined maximum length $L$. Each discovered path is converted into its corresponding relational path, which forms a candidate rule body $\mathbf{r}_{b}$. After enumerating all candidate rules, we quantify their quality using established statistical metrics. For each rule $r_{h}\leftarrow\mathbf{r}_{b}$, we compute its standard confidence, defined as the conditional probability $P(r_{h}|\mathbf{r}_{b})$,
@@ 4.1.2 Query-Centric Subgraph Extraction
Feature engineering. A raw subgraph is insufficient for a GNN to distinguish node roles. We therefore enrich the subgraph with node features $\mathbf{x}_{v}$. Following SEAL (Zhang & Chen, 2018) and GraIL (Teru et al., 2020), we employ relative node labeling to capture structural patterns independent of node identities. The feature vector $\mathbf{x}_{v}$ is a concatenation of three components encoding the node’s structural role relative to the query entity $h$.
(i) Head Entity Indicator: A two-dimensional binary vector that identifies whether a node is the query head entity $h$ (i.e., $[1,0]$) or a neighbor (i.e., $[0,1]$). This anchors the subgraph representation around the query. (ii) Shortest Path Distance: The geodesic distance (i.e., minimum number of hops) from node $v$ to the head entity $h$ within the subgraph $\mathcal{G}_{h}$. This explicitly encodes the structural proximity of each node to the query’s origin. (iii) Global Centrality Score: The log-scaled degree of node $v$ as calculated from the complete graph $\mathcal{G}$. This feature injects a measure of the node’s global importance into the local subgraph context.
@@ 4.1.3 Instance creation and Negative Sampling
For a query $(h_{i},r_{i},t_{i})$, let ${\bf R}_{i}^{+}$ and ${\bf R}_{i}^{-}$ consist of the $k_{pos}$ positive rules $r_{i}\leftarrow r_{b_{i1}}^{+},\cdots,r_{i}\leftarrow r_{b_{ik_{pos}}}^{+}$} and $k_{neg}$ hard negative rules $r_{i}\leftarrow r_{b_{i1}}^{-},\cdots,r_{i}^{-}\leftarrow r_{b_{ik_{neg}}}^{-}$} respectively. The instances generated in this step, subsequently provided to our model, comprise rule-enriched triplets, $((h_{1},r_{1},t_{1}),{\bf R}_{1}^{+}$, ${{\bf R}_{1}^{-}})$, $((h_{2},r_{2},t_{2}),{\bf R}_{2}^{+}$, ${\bf R}_{2}^{-})$, etc.
@@ 4.3 Inference and Final Ranking
Input: query $\mathbf{q}(h,r,?)$, model $\phi(\cdot)$, hyperparameters Output: a vector of scores that measure the plausibility of different entities answering $\mathbf{q}$ 1. Candidate rule generation from global rule base 2. For each rule $r\leftarrow\mathbf{r}_{b}$, predict $\phi(h,r,\mathbf{r}_{b})$ from model $\phi(\cdot)$ 3. Compute adjusted scores $\phi^{\prime}$ using penalty $\lambda$, then update to confidence $w_{i}$ using softmax with temperature $T$. 4. For all the entities, calculate their potential to answer $\mathbf{q}$ (a) Calculate a rule grounding score $ground(h,\mathbf{r}_{b})$ (b) Apply a binarization function to the grounding score $B(ground(h,\mathbf{r}_{b}))$ (c) Compute $\mathbf{v}_{\text{ans}}$ using equation 2, which represents the entities potential to answer $\mathbf{q}$ 5. Return $\mathbf{v}_{\text{ans}}$
@@
Figure 2 shows the steps to answer the given query. Step 1 generates candidate rules. For any given query $(h,r,?)$, exhaustively scoring all rules in the global rule base would be computationally prohibitive. To obtain a manageable yet high-quality candidate set, we first identify all rules that are locally applicable to $h$, and then select the top-$N$ among them based on their Wilson score.
Step 2 calculates a context-aware (localized) importance score for each candidate rule. In particular, each candidate rule $\mathbf{r}_{b_{i}}$ is passed to our trained scoring function $\phi(h,r,\mathbf{r}_{b_{i}})$, which leverages the query’s unique subgraph context $\mathcal{G}_{h}$. This step assigns a dynamic, context-aware score to each rule, moving beyond static metrics to assess its true relevance for the specific query. To mitigate the influence of overly broad rules which often introduce ambiguity in dense graphs, we calculate an adjusted score $\phi^{\prime}(h,r,\mathbf{r}_{b})=\phi(h,r,\mathbf{r}_{b})-\lambda\log(n_{tails})$, where $n_{tails}$ represents the number of distinct tail entities reachable by the rule and $\lambda$ is a penalty coefficient (refer to Sec. 5.2 for analysis). Next, the context-aware scores are converted into a confidence distribution using a temperature-controlled softmax (Step 3). The weight $w_{i}$ for each rule is calculated as $w_{i}=\frac{\exp(\phi^{\prime}_{i}/T)}{\sum_{j}\exp(\phi^{\prime}_{j}/T)}$. The temperature $T$ is an important hyperparameter. As $T\to 0$, the softmax approximates an argmax function, concentrating all weight onto the single highest-scoring rule. This allows the model to rely solely on the strongest contextual signal.
@@
A naive grounding score can be a noisy signal dominated by high-degree nodes. To mitigate this effect, we squash the path counts using the tanh function, controlled by a tanh scale hyperparameter $\tau$, $B(\mathbf{g}_{i})=\tanh(\mathbf{g}_{i}/\tau)$. This provides a soft, saturated count that is controlled by the tanh scale hyperparameter $\tau$, preventing an unbounded influence from numerous paths.
Finally (Step 4c), a vector $\mathbf{v}_{\text{ans}}$ that captures all the entities’ plausibility to answer the given query is computed as the weighted sum of all the binarized, grounded rule vectors as in equation 2.
@@ 5 Experiments: SLogic’s performance in KGC
Further details on the experimental setting, including default setting of hyperparameters, hardware configuration, implementation specifics, and training procedures, are provided in Appendix B.